# Quantum field theory treatment of the neutrino spin-flavor precession in a magnetic field

### v2

@@ 1 Introduction
As a rule, neutrino oscillations in vacuum and in external fields are described within the quantum mechanical approach, which gives satisfactory results in the majority of situations. Nonetheless, this formalism has some shortcomings which were outlined in Ref. [12]. Numerous attempts to construct the description of neutrino flavor oscillations based on the quantum field theory (QFT) methods are reviewed in Ref. [13].
@@
This work is organized in the following way. We remind briefly some basic issues of the Majorana neutrinos electrodynamics in Sec. 2. Then, in Sec. 3, we outline the QFT approach and adapt it for the neutrino spin-flavor precession. We derive the dressed propagators for the neutrino mass eigenstates in a magnetic field by solving the corresponding Dyson equations in Sec. 4. The matrix element and the transition probability for the neutrino spin-flavor precession are computed in Sec. 5. The results of the QFT and the standard quantum mechanical approaches are compared in Sec. 6. Finally, we discuss our results in Sec. 7. The propagators of massive Weyl neutrinos in vacuum are derived in Appendix.
@@ 3 Quantum field theory formalism for neutrino oscillations
Following Ref. [14], we assume that the source and the detector of neutrinos, separated by the distance $\mathbf{L}$, consist of heavy nuclei $N$ and $N^{\prime}$. The production and detection of neutrinos is characterized by the interaction of these nuclei with charged leptons $l_{\alpha}^{-}$ and antileptons $l_{\beta}^{+}$, where $\alpha,\beta=e,\mu,\dotsc$. Thus, neutrino oscillations correspond to the process $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$. If the flavor content of charged leptons changes, i.e. $\beta\neq\alpha$, one can speak about neutrino flavor oscillations. If electric charges of leptons also change, e.g., we have an electron in a source and antimuon in a detector, such a process is allowed in vacuum for Majorana neutrinos. It was connected to neutrino-to-antineutrino oscillations in Refs. [14, 24]. However, if we study the transition $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$ in an external field which neutrinos interact with, it is more appropriate to call it the neutrino spin-flavor precession. The considered process is schematically depicted in Fig. 1.
Figure 1: The schematic Feynman diagram for the reaction $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$. The neutrino line has two counter-directional arrows since we deal with the Majorana particles. Moreover, the neutrino line is broad since the mass eigenstates interact with a magnetic field [see Eq. (2)]. Note that we take into account the magnetic interaction exactly (see Sec. 4).
@@
We assume that the nuclei in the source and the detector are quite heavy. Hence, their positions, $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, are unchanged, with $\mathbf{x}_{2}-\mathbf{x}_{1}=\mathbf{L}$. In this case, the nuclear currents in Eq. (3) have the form,
@@
$S=4\pi G_{\mathrm{int}}^{2}e^{-i\mathbf{p}_{\alpha}\mathbf{x}_{2}+i\mathbf{p}_{\beta}\mathbf{x}_{1}}\delta(E_{\alpha}-E_{\beta})\mathcal{M}_{\beta\to\bar{\alpha}},$ (3.8)
where the matrix element is
@@ 4 Propagators of massive Majorana neutrinos in a magnetic field
Figure 2: Feynman diagrams corresponding to Eq. (4). 4D Fourier images of the bare propagators $\mathrm{i}S_{a\mathrm{R,L}}$, given in Eqs. (A.2) and (A.10), are depicted by thin solid lines [25, pp. 468–472]. The broad solid lines correspond to the dressed propagators $\mathrm{i}\Sigma_{ab}$. The potential of the magnetic interaction is $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$.
@@ 5 Transition probability of the spin-flavor precession
We present the detailed calculation of the $\Sigma_{12}$ contribution to the matrix element. The magnetic field is taken to be transverse with respect $\mathbf{L}$, $\mathbf{B}\perp\mathbf{L}$. For instance, we suppose that $\mathbf{L}=L\mathbf{e}_{z}$ and $\mathbf{B}=B\mathbf{e}_{x}$, where $\mathbf{e}_{x,z}$ are the unit vectors along the $x$- and $z$-axes. In the quantum mechanical approach, a neutrino is on the mass shell, i.e. the neutrino momentum is also perpendicular to $\mathbf{B}$. In the QFT treatment, a neutrino is a virtual particle. Thus, $\mathbf{q}$ is arbitrary and, hence, $(\hat{q}\cdot\mathbf{B})\neq 0$ in Eq. (4.5).
@@
$z_{+}$ $=\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}\times\begin{cases}\sqrt{\rho_{0}^{2}-\rho^{2}},&\text{if}\ \rho<\rho_{0},\\ \mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}},&\text{if}\ \rho>\rho_{0},\end{cases}$ (5.8)
@@
$P_{\mathrm{max}}^{(\mathrm{qm})}=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}},\quad P_{\mathrm{max}}^{(\mathrm{qft})}=-\frac{(\mu B)^{4}}{E\left[(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}\right]^{3/2}}.$ (5.16)
We shall see shortly in Sec. 6 that $P_{\mathrm{max}}^{(\mathrm{qm})}$ in Eq. (5.16) reproduces the result of the quantum mechanical approach. The term $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16) is the QFT contribution to the transition probability. It arises from the terms (i) and (ii) in the propagator $\Sigma_{12}$
@@
One can see in Eq. (5.16) that the QFT contribution to the transition probability is negative. Moreover, $|P_{\mathrm{max}}^{(\mathrm{qft})}|/P_{\mathrm{max}}^{(\mathrm{qm})}\sim\mu B/E\sim\mathfrak{E}/E$. Thus, the QFT term is much smaller than the quantum mechanical one for all realistic oscillations channels where neutrinos are ultrarelativistic.
The latter result proves the assumption that the terms (i) and (ii) in $\Sigma_{12}$ in Eq. (4.5) can be safely neglected for ultrarelativistic neutrinos since their contribution to the transition probability is small. We recall that all the terms in $\Sigma_{12}$ in Eq. (4.5), including (i) and (ii), are of the same order of magnitude. Therefore, before the transition probability computation, one cannot say whether certain terms in $\Sigma_{12}$ are smaller or greater.
@@ 5.1 Scattering of charged leptons with arbitrary energies
i.e. we remove the QFT corrections from Eq. (5.14) and take into account the additional factor $e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\tfrac{\theta_{\bar{\mu}}}{2}\right)\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}/2$.
The measurable quantity of the process $e+N\to\bar{\mu}+\tilde{N}^{\prime}$ is the differential cross-section [25, p. 252],
@@
One can see in Eq. (5.21) that the total cross-section oscillates in space and is proportional to the quantum mechanical transition probability, $P_{\mathrm{max}}^{(\mathrm{qm})}\sin^{2}(\mathfrak{E}L)$; cf. Eqs. (5.15) and (5.16).
@@ 6 Quantum mechanical description of the spin-flavor precession
In this section, we analyze the quantum mechanical description of the spin-flavor precession of Majorana neutrinos in an external magnetic field. This study can be done in both flavor and mass bases. We take the flavor basis since it is more common in the standard quantum mechanical approach.
@@
Now, we can compare the quantum mechanical and quantum field theory approaches to the neutrino spin-flavor precession. If neutrinos are ultrareletivistic, we can put $E_{\nu}=E$, $\mathfrak{E}_{\nu}=\mathfrak{E}$, and $t\approx L$ in Eq. (6.23). Then, we omit the small correction $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16). Under these approximations, the transition probability in Eqs. (5.15) and (5.16) coincides with that in Eq. (6.23), i.e. both approaches provide the same description of the spin-flavor precession.
@@ 7 Discussion
In the present work, we have developed the QFT description of neutrino the spin-flavor precession in a magnetic field. We were based on the QFT formalism, which was originally developed in Refs. [14, 15, 16], where neutrinos are virtual particles. This method was used in Ref. [14] to study neutrino-to-antineutrino oscillations in vacuum. If one deals with Majorana neutrinos with transition magnetic moments, $\nu\leftrightarrow\bar{\nu}$ oscillations are analogous to spin-flavor transitions in a magnetic field. We have adapted the results of Ref. [14] for the description of the spin-flavor precession.
@@
Finally, we have derived the transition probability for the $\nu_{e}\to\bar{\nu}_{\mu}$ precession in a magnetic field in Sec. 5 within the QFT approach. This transition probability has the leading term, which was shown in Sec. 6 to coincide with the result of the standard quantum mechanical approach. Moreover, we have obtained that the dressed propagator in Eq. (4.5) acquires the terms (i) and (ii) which have the QFT nature. These terms cannot be neglected before the transition probability calculation. We have demonstrated in Sec. 5 that the QFT terms in the propagator result in the correction to the leading term in the transition probability which is small if neutrinos are ultrarelativistic; cf. Eqs. (5.15) and (5.16).
@@
The most important approximation is the ultrarelativity of virtual neutrinos. It can be expressed in the form, $E=(E_{\alpha}+E_{\beta})/2\gg m_{1,2}$. Along with the spin-flavor precession under the influence of a magnetic field, for Majorana particles, one has neutrino-to-antineutrino transitions owing to a nonzero neutrino mass. Thus, the series in Eq. (4), would have a more complicated form. Moreover, the bare propagators, derived in Appendix, have rather simple form for ultrarelativistic neutrinos. Finally, without the assumption of the neutrinos ultrarelativity, the computation of the contour integrals, $I_{+}$ and $I_{-}$, would be impossible in Sec. 5. For neutrinos with arbitrary energies, the equation $E-\mathcal{E}_{\pm}(\mathbf{q})+\mathrm{i}0=0$, which defines the poles, has eight roots which should found analytically. Thus, the assumption of ultrarelativistic neutrinos in crucial to treat the spin-flavor precession in frames of QFT.
@@
In deriving of Eq. (5), we made an assumption of the great propagation distance $L$. It allows one to avoid terms $\propto L^{-2}$ in Eq. (5). It is, however, known that the coherence in neutrino oscillations is lost if $L>L_{\mathrm{coh}}$, where $L_{\mathrm{coh}}$ is the coherence length. It is found that $L_{\mathrm{coh}}\propto L_{\mathrm{osc}}\sim\mathfrak{E}^{-1}$ (see, e.g., Ref. [12]).
@@
In our case, when we study neutrinos interacting with an external magnetic field, the integrations in Sec. 5 are quite nontrivial even for plane waves charged leptons. As we mentioned earlier, the matrix element can be derived only in the approximation of ultrarelativistic neutrinos. Therefore, a careful consideration of the decoherence in the spin-flavor precession in frames of QFT is a quite complicated task which is beyond the scope of the present work. We can just roughly estimate the coherence length as $L_{\mathrm{coh}}\sim\sigma_{x}E/\mathfrak{E}$ for ultrarelativistic neutrinos [15]. Another estimate for the coherence length in a magnetic field reads [27] $L_{\mathrm{coh}}\sim\sigma_{x}E^{3}/\mathfrak{E}m_{a}^{2}$.

### v3

@@ 1 Introduction
As a rule, neutrino oscillations in vacuum and in external fields are described within the quantum mechanical (QM) approach, which gives satisfactory results in the majority of situations. The QM formalism is based on the following assumptions. Flavor neutrinos are the coherent superposition of mass states. All mass states have definite equal momenta while evolving in time. The mass states are ultrarelativistic. The propagation time of a neutrino beam is approximately equal to the propagation distance. Nonetheless, these assumptions were shown in Ref. [12] to require the justification based upon a quantum field theory (QFT) formalism. Moreover, the effective wavefunction, which one deals with in QM approach, has to be somehow related to the true spinor wavefunctions of neutrinos. We should admit that the necessity of the QFT application to describe neutrino oscillations was realized long time ago. Numerous attempts to formulate a QFT based approach to neutrino oscillations are reviewed in Ref. [13].
@@
This work is organized in the following way. We remind briefly some basic issues of the Majorana neutrinos electrodynamics in Sec. 2. Then, in Sec. 3, we outline the QFT approach and adapt it for the neutrino spin-flavor precession. We derive the dressed propagators for the neutrino mass eigenstates in a magnetic field by solving the corresponding Dyson equations in Sec. 4. The matrix element and the transition probability for the neutrino spin-flavor precession are computed in Sec. 5. The results of the QFT and the standard QM approaches are compared in Sec. 6. Finally, we discuss our results in Sec. 7. The propagators of massive Weyl neutrinos in vacuum are derived in Appendix.
@@ 3 Quantum field theory formalism for neutrino oscillations
Following Ref. [14], we assume that the source and the detector of neutrinos, separated by the distance $\mathbf{L}$, consist of heavy nuclei $N$ and $N^{\prime}$ at $t\to$-$\infty$. In addition to these nuclei, we have the charged lepton $l_{\beta}^{-}$, where $\beta=e,\mu,\dotsc$, in the in-state. Since the charge lepton has a definite mass, its in-state is well defined. Then, we take that $l_{\beta}^{-}$ interacts with the source, i.e. with the nucleus $N$. Neutrino states are produced in this interaction: $l_{\beta}^{-}+N\to\tilde{N}+\text{neutrinos}$. Note that a nuclear transformation takes place in a source, i.e. $N\neq\tilde{N}$, since electric charges of $l_{\beta}^{-}$ and neutrinos are different. The interaction of $l_{\beta}^{-}$ with the source corresponds to the left hand vertex in Fig. 1.
Then, the neutrino beam travels towards a detector consisting of a nucleus $N^{\prime}$. While reaching it, neutrinos are absorbed by $N^{\prime}$: $\text{neutrinos}+N^{\prime}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$, where $l_{\alpha}^{+}$ is the anti-lepton of the flavor $\alpha=e,\mu,\dotsc$ and $\tilde{N}^{\prime}$ is the new nucleus different from $N^{\prime}$. Therefore, at $t\to+\infty$ one has $\tilde{N}$ in a source, $\tilde{N}^{\prime}$ in a detector, as well as $l_{\alpha}^{+}$. Analogously to $l_{\beta}^{-}$, the anti-lepton $l_{\alpha}^{+}$ also has a definite mass. Thus, the out-state of $l_{\alpha}^{+}$ at $t\to+\infty$ is also well defined. The neutrino interaction with a source corresponds to the vertex in the right hand side in Fig. 1.
Now, we suppose that the flavor content of charged leptons changes, i.e. $\beta\neq\alpha$. In this situation, one can speak about neutrino flavor oscillations. Moreover, we assume that electric charges of leptons are opposite. For example, we have an electron in a source and antimuon in a detector. Such a process is allowed in vacuum only for Majorana neutrinos. It was connected to neutrino-to-antineutrino oscillations in Refs. [14, 24]. That is why we place counter-directional arrows over the neutrino line in Fig. 1 to emphasize that the lepton number is violated. However, if we study the transition $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$ in an external field which neutrinos interact with, it is more appropriate to call it the neutrino spin-flavor precession.
Figure 1: The schematic Feynman diagram for the reaction $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$. The arrows over fermionic lines, $l_{\beta}^{-}$, $l_{\alpha}^{+}$, and neutrinos, show the flow of the lepton number. Since we do not care about the details of the nuclear transformations in the source and the detector, the arrows over nuclear lines, NN, $\tilde{N}$, $N^{\prime}$, and $\tilde{N}^{\prime}$, can signify the baryon number flow. The neutrino line has two counter-directional arrows since we deal with the Majorana particles which violate the lepton number. The time flows from left to right. Moreover, the neutrino line is broad since the mass eigenstates interact with a magnetic field [see Eq. (2)]. Note that we take into account the magnetic interaction exactly (see Sec. 4).
@@
We choose the simplest model of the source and the detector. We assume that the nuclei $N$ and $N^{\prime}$, as well as $\tilde{N}$ and $\tilde{N}^{\prime}$, are quite heavy. Hence, their positions, $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, are unchanged, with $\mathbf{x}_{2}-\mathbf{x}_{1}=\mathbf{L}$. In this case, the nuclear currents in Eq. (3) have the form,
@@
i.e. the nuclear matrix elements are time independent. Therefore, the consideration of the source and the detector on the classical level allows one to factorize the nuclear contribution in the $S$-matrix in Eq. (3).
@@
$S=4\pi G_{\mathrm{int}}^{2}e^{-i\mathbf{p}_{\alpha}\mathbf{x}_{2}+i\mathbf{p}_{\beta}\mathbf{x}_{1}}\delta(E_{\alpha}-E_{\beta})\mathcal{M}_{\beta\to\bar{\alpha}}.$ (3.8)
One can see in Eq. (3.8) that the total energy is conserved, $E_{\alpha}=E_{\beta}$. This fact results from the assumption that nuclei in the source and the detector are static. Moreover, formally, we consider the situation like a scattering, i.e. all external particles are in- and out-states at $t\to\mp\infty$. Hence, one has the infinite observation time $\infty<t<+\infty$. It allows us to integrate over the infinite time intervals in Eq. (3) to get the energy conservation delta-function in Eq. (3.8). The violation of the energy conservation assumption would lead to the decoherence in neutrino oscillation, which is discussed shortly in Sec. 7.
The matrix element in Eq. (3.8) is
@@
Earlier, we mentioned that the marker of the spin-precession is the appearance of the anti-lepton $l_{\alpha}^{+}$ (out-state) in the detector provided one had the lepton $l_{\beta}^{-}$ (in-state) in the source. It leads to the bi-spinors $v_{\alpha}$ and $u_{\beta}$, as well as the propagator of massive neutrinos $\propto\left\langle 0\left|T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\right|0\right\rangle$ in Eq. (3). Note that this propagator is nonzero only when neutrinos are Majorana particles. That is why, virtual neutrinos undergo the transitions $\nu_{\beta}\to\bar{\nu}_{\alpha}$ effectively. Since only left neutrinos are active, we may say that we deal with the transitions $\nu_{\beta\mathrm{L}}\to\bar{\nu}_{\alpha\mathrm{L}}$.
If we studied the analogue of the process in Fig. 1 for Dirac neutrinos, one would have leptons $l_{\alpha,\beta}^{-}$ with the same electric charge interacting with a detector and a source (out- and in-states). Therefore, in Eq. (3), we would have the bi-spinors $u_{\alpha,\beta}$, as well as the propagator $\propto\left\langle 0\left|T\left\{\psi_{a}(x)\bar{\psi}_{b}(y)\right\}\right|0\right\rangle$ for Dirac neutrinos. This situation would correspond to the effective transformation of virtual neutrinos like $\nu_{\beta}\to\nu_{\alpha}$ or $\nu_{\beta\mathrm{L}}\to\nu_{\alpha\mathrm{L}}$.
Finally, we mention that despite the similarity in the description of spin-flavor precession of Majorana and Dirac neutrinos (see, e.g., Refs. [7, 9]), there is a conceptual difference between these two processes. The spin-flavor precession of Dirac particles allows a quasiclassical description since one can introduce an effective spin vector which precesses in an external magnetic field [25]. For Majorana neutrinos, it is the essentially quantum process since it involves transitions between particles and antiparticles.
@@ 4 Propagators of massive Majorana neutrinos in a magnetic field
Figure 2: Feynman diagrams corresponding to Eq. (4). 4D Fourier images of the bare propagators $\mathrm{i}S_{a\mathrm{R,L}}$, given in Eqs. (A.2) and (A.10), are depicted by thin solid lines [26, pp. 468–472]. The broad solid lines correspond to the dressed propagators $\mathrm{i}\Sigma_{ab}$. The potential of the magnetic interaction is $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$.
@@ 5 Transition probability of the spin-flavor precession
We present the detailed calculation of the $\Sigma_{12}$ contribution to the matrix element. The magnetic field is taken to be transverse with respect to $\mathbf{L}$, $\mathbf{B}\perp\mathbf{L}$. For instance, we suppose that $\mathbf{L}=L\mathbf{e}_{z}$ and $\mathbf{B}=B\mathbf{e}_{x}$, where $\mathbf{e}_{x,z}$ are the unit vectors along the $x$- and $z$-axes. In the QM approach, a neutrino is on the mass shell, i.e. the neutrino momentum is also perpendicular to $\mathbf{B}$. In the QFT treatment, a neutrino is a virtual particle. Thus, $\mathbf{q}$ is arbitrary and, hence, $(\hat{q}\cdot\mathbf{B})\neq 0$ in Eq. (4.5).
@@
$z_{+}$ $=\sqrt{1+\frac{(\mu B)^{2}}{E\mathfrak{E}}\cos^{2}\phi}\times\begin{cases}\sqrt{\rho_{0}^{2}-\rho^{2}}+\mathrm{i}0,&\text{if}\ \rho<\rho_{0},\\ \mathrm{i}\sqrt{\rho^{2}-\rho_{0}^{2}},&\text{if}\ \rho>\rho_{0},\end{cases}$ (5.8)
@@
$P_{\mathrm{max}}^{(\mathrm{qm})}=\frac{(\mu B)^{2}}{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}},\quad P_{\mathrm{max}}^{(\mathrm{qft})}=-P_{\mathrm{max}}^{(\mathrm{qm})}\frac{(\mu B)^{2}}{E\sqrt{(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}}}.$ (5.16)
We shall see shortly in Sec. 6 that $P_{\mathrm{max}}^{(\mathrm{qm})}$ in Eq. (5.16) reproduces the result of the QM approach. The term $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16) is the QFT contribution to the transition probability. It arises from the terms (i) and (ii) in the propagator $\Sigma_{12}$
@@
One can see in Eq. (5.16) that the QFT contribution to the transition probability is negative. Moreover, $|P_{\mathrm{max}}^{(\mathrm{qft})}|/P_{\mathrm{max}}^{(\mathrm{qm})}\sim\mu B/E\sim\mathfrak{E}/E$. Thus, the QFT term is much smaller than the QM one for all realistic oscillations channels where neutrinos are ultrarelativistic.
The latter result proves the assumption that the terms (i) and (ii) in $\Sigma_{12}$ in Eq. (4.5) can be safely neglected for ultrarelativistic neutrinos since their contribution to the transition probability is small. We recall that all the terms in $\Sigma_{12}$ in Eq. (4.5), including (i) and (ii), are of the same order of magnitude. Therefore, before the transition probability computation, one cannot say anything about the magnitude of certain terms in $\Sigma_{12}$.
@@ 5.1 Scattering of charged leptons with arbitrary energies
i.e. we remove the QFT corrections from Eq. (5.14) and take into account the additional factor, $e^{\mathrm{i}\phi_{\bar{\mu}}}\cos\left(\tfrac{\theta_{\bar{\mu}}}{2}\right)\sqrt{(1+v_{e})(1+v_{\bar{\mu}})}/2$.
The measurable quantity of the process $e+N\to\bar{\mu}+\tilde{N}^{\prime}$ is the differential cross-section [26, p. 252],
@@
One can see in Eq. (5.21) that the total cross-section oscillates in space and is proportional to the QM transition probability, $P_{\mathrm{max}}^{(\mathrm{qm})}\sin^{2}(\mathfrak{E}L)$; cf. Eqs. (5.15) and (5.16).
@@ 6 Quantum mechanical description of the spin-flavor precession
In this section, we analyze the QM description of the spin-flavor precession of Majorana neutrinos in an external magnetic field. This study can be done in both flavor and mass bases. We take the flavor basis since it is more common in the standard QM approach.
@@
Now, we can compare the QM and QFT approaches to the neutrino spin-flavor precession. If neutrinos are ultrareletivistic, we can put $E_{\nu}=E$, $\mathfrak{E}_{\nu}=\mathfrak{E}$, and $t\approx L$ in Eq. (6.23). Then, we omit the small correction $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16). Under these approximations, the transition probability in Eqs. (5.15) and (5.16) coincides with that in Eq. (6.23), i.e. both approaches provide the same description of the spin-flavor precession.
@@ 7 Discussion
In the present work, we have developed the QFT description of the neutrino spin-flavor precession in a magnetic field. We were based on the QFT formalism, which was originally developed in Refs. [14, 15, 16], where neutrinos are virtual particles. This method was used in Ref. [14] to study neutrino-to-antineutrino oscillations in vacuum. If one deals with Majorana neutrinos with transition magnetic moments, $\nu\leftrightarrow\bar{\nu}$ oscillations are analogous to spin-flavor transitions in a magnetic field. We have adapted the results of Ref. [14] for the description of the spin-flavor precession.
@@
Finally, we have derived the transition probability for the $\nu_{e}\to\bar{\nu}_{\mu}$ precession in a magnetic field in Sec. 5 within the QFT approach. This transition probability has the leading term, which was shown in Sec. 6 to coincide with the result of the standard QM approach. The coincidence of the predictions of QFT and QM approaches takes place for ultrarelativistic neutrinos. In this situation, we identify the distance between the source and the detector with the propagation time, $L\approx t$. Additionally, one can neglect the QFT contribution to the transition probability for ultrarelativistic neutrinos.
Moreover, we have obtained that the dressed propagator in Eq. (4.5) acquires the terms (i) and (ii) which have the QFT nature. These terms cannot be neglected before the transition probability calculation. We have demonstrated in Sec. 5 that the QFT terms in the propagator result in the correction to the leading term in the transition probability which is small if neutrinos are ultrarelativistic; cf. Eqs. (5.15) and (5.16).
Despite we treat neutrinos as virtual particles, the consideration of the QFT correction to the transition probability shows that the virtuality of neutrinos is negligible. Indeed, in Sec. 5, we have demonstrated that the terms (i) and (ii) in $\Sigma_{12}$ can be omitted in a realistic situation. Hence, virtual neutrinos behave as if they were on the mass shell and were described within the QM formalism. In particular, it happens since we study macroscopically large propagation distances.
@@
The comprehensive application of QFT to the description of neutrino oscillations and, in particular, to the spin-flavor precession, should involve various loop effects along with the diagrams where neutrinos participate on a tree level. We have accounted for some of these loop contributions indirectly. For example, we have assumed that the neutrino magnetic moment already has a certain value instead of calculating the neutrino electromagnetic vertex function (see, e.g., Ref. [27]) at any interaction of the undressed neutrino propagators $S_{a\mathrm{L,R}}$ with the external magnetic potentials $\pm\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$ in Fig. 2. The same argument is addressed to the neutrino masses $m_{1,2}$ which are supposed to have physical values. Here, we assume that the model underlying the neutrino mass generation is renormalizable.
Another group of vacuum contributions to neutrino oscillations appears in the approach [28], where the Fock space of neutrino flavor eigenstates was constructed. It was obtained that a rapidly oscillating term in the transition probability arises along with the conventional slowly varying Pontecorvo contribution [29, 30]. This new contribution, resulting from the nontrivial flavor vacuum structure, was shown to vanish for utraretivistic neutrinos. Note that analogous point of view to neutrino oscillations was suggested in Ref. [31].
The formalism developed in Refs. [28, 29, 30] describes transitions between different neutrino flavors happening in time. In practice, this kind of transitions can be implemented if one has a permanent source of neutrinos (see also Ref. [21]). Note that the diagrammatic approach to the evolution of the interacting fields in a finite time interval was elaborated in Ref. [32].
Additionally, the approach, where the mixing between different neutrino types was considered as the perturbation, was developed in Ref. [30]. It is interesting to mention that the mixing was accounted for in Ref. [30] by solving the Dyson equation (see also Ref. [33]) analogously to our approach where we use the Dyson like equations to take into account the external magnetic field.
In our work, we follow the QFT approach to neutrino oscillations based on the conventional Fock states of the neutrino mass eigenstates [14, 15, 16] (see, e.g., Eq. (3.10) where the dressed propagators are defined). The advantage of this formalism consists in the fact one does not need to bother about the Poincaré symmetry properties of the Fock states, including the vacuum one; cf. Refs. [30, 31]. We also mention that our analysis does not reveal rapidly oscillating vacuum terms in the transition probability predicted in Refs. [29, 30].
@@
The most important approximation is the ultrarelativity of virtual neutrinos. It can be expressed in the form, $E=(E_{\alpha}+E_{\beta})/2\gg m_{1,2}$. Along with the spin-flavor precession under the influence of a magnetic field, for Majorana particles, one has neutrino-to-antineutrino transitions owing to a nonzero neutrino mass. Thus, the series in Eq. (4), would have a more complicated form. Moreover, the bare propagators, derived in Appendix, have rather simple form for ultrarelativistic neutrinos. Finally, without the assumption of the neutrinos ultrarelativity, the computation of the contour integrals, $I_{+}$ and $I_{-}$, would be impossible in Sec. 5. For neutrinos with arbitrary energies, the equation $E-\mathcal{E}_{\pm}(\mathbf{q})+\mathrm{i}0=0$, which defines the poles, has eight roots which should be found analytically. Thus, the assumption of ultrarelativistic neutrinos in crucial to treat the spin-flavor precession in frames of QFT.
@@
In deriving of Eq. (5), we made an assumption of the great propagation distance $L$. It allows one to avoid terms $\propto L^{-2}$ in Eq. (5). It is, however, known that the coherence in neutrino oscillations is lost if $L>L_{\mathrm{coh}}$, where $L_{\mathrm{coh}}$ is the coherence length. The decoherence effects in neutrino oscillations in vacuum in frames of both QM and QFT approaches have been studied in Refs. [12, 34].
@@
In our case, when we study neutrinos interacting with an external magnetic field, the integrations in Sec. 5 are quite nontrivial even for plane waves charged leptons. As we mentioned earlier, the matrix element can be derived only in the approximation of ultrarelativistic neutrinos. Therefore, a careful consideration of the decoherence in the spin-flavor precession in frames of QFT is a quite complicated task which is beyond the scope of the present work.
We can only say that the coherence length is infinite in our case [15]. It results from the fact that we use the plane wave approximation for charged leptons wavefunctions in Eq. (3.3), as well as from the assumption of the exact energy conservation in Eq. (3.8). As we mentioned above, the plane waves and the energy conservation approximations are desirable but not crucial for our analysis. Nevertheless, the decoherence effects have not been studied in our work since we have adopted these approximations to advance in analytical calculations in Sec. 5.